OpenStudy (anonymous):

write equation in slope-intercept form for the line passing through points (-3,7) and (2,-3)

OpenStudy (anonymous):

First you find the slope, which is -2.

OpenStudy (anonymous):

Then you find the y intercept which I think is -4

OpenStudy (anonymous):

But I'm not completely sure someone please back me up one that or check me.

OpenStudy (anonymous):

If -4 is the y intercept then this equation in slope intercept form is y=-2x-4

OpenStudy (anonymous):

\[(-3;7) and (2;-3)\] \[m = \frac{ -3-7 }{ 2-(-3) }=-2\] \[y=mx+c\] \[using the point (-3;7)\] \[7=-2(-3)+c\] \[7=6+c \therefore c=1\] \[y=mx+c\] \[y=-2x+1\]